# Splittings of One-Ended Groups with One-Ended Halfspaces

### v3

@@ 4.2 Chopping up a halfspace $\mathfrak{h}_{0}$
[…]
Figure 1: Cartoon picture for the proof of Lemma 4.9.
@@
The modification of $X$ we consider is where we fix an integer $n\geq 1$, and replace each edge in each wall of $X$ with $n$ edges that join the same pair of vertices. Let $X_{n}$ be the resulting graph, which is endowed with a collection of halfspaces and walls that are inherited from $X$ in the obvious way. Any halfspace cut $(\mathfrak{h},C)$ in $X$ is still a halfspace cut in $X_{n}$, and vice versa. The only difference is that the edge boundary $\delta C$ might become larger in $X_{n}$ because of the extra edges. More precisely, the numbers $W(C)$ and $|\delta C|$ both increase by $(n-1)W(C)$ when passing from $X$ to $X_{n}$. Note that a halfspace cut is $W$-minimal with respect to $X$ if and only if it is $W$-minimal with respect to $X_{n}$, but the property of having minimal $|\delta C|$ may change when passing between $X$ and $X_{n}$. Let $w$ be the smallest value of $W(C)$ among halfspace cuts $(\mathfrak{h},C)$ in $X$, and let $N$ be the smallest value of $|\delta C|$ among $W$-minimal halfspace cuts $(\mathfrak{h},C)$ in $X$. Put $n=N+1$. Then the smallest possible value of $|\delta C|$ in $X_{n}$ is $N+(n-1)w$, and this is only achieved by $W$-minimal halfspace cuts, as required. ∎
Now let $(\mathfrak{h}_{0},C)$ be a halfspace cut with $|\delta C|$ minimal. Lemmas 4.9 and 4.10 imply that each halfspace $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$ satisfies either $V\mathfrak{h}\subseteq C$ or $V\mathfrak{h}\subseteq V\mathfrak{h}_{0}-C$. Furthermore, by Theorem 4.5 we can assume that the pocset $\mathcal{P}(\mathfrak{h}_{0},C)$ has no pairs of transverse elements.
@@ 4.3 The trees $T_{e}$
For brevity we will write $G_{0}=G_{\mathfrak{h}_{0}}$ for the $G$-stabilizer of $\mathfrak{h}_{0}$ and we will write $\delta$ for the edge boundary $\delta C$. As in Subsection 4.1, we can form the pocset $\mathcal{P}(\mathfrak{h}_{0},C)$. Now consider the subpocset
@@
We can then define the trees $T_{e}$ required for Theorem 4.1, and the action of $G$ on $\sqcup_{e}T_{e}$. Indeed if $e_{0}\in ET$ is the edge corresponding to $\mathfrak{h}_{0}$ (i.e. $p(\mathfrak{h}_{0})\subseteq T$ is bounded by $\hat{e}_{0}$), then put
@@
for each $i\in\Omega$. Note that the $G$-stabilizer of $T_{e_{0}}$ is $G_{0}=G_{e_{0}}$, and the action $G_{0}\curvearrowright T_{e_{0}}$ is a non-trivial finite splitting with finite edge stabilizers by Lemmas 4.11 and 4.12. Similarly, the $G$-stabilizer of $T_{g_{i}e_{0}}$ is $g_{i}G_{0}g_{i}^{-1}=G_{g_{i}e_{0}}$, and the action $G_{g_{i}e_{0}}\curvearrowright T_{g_{i}e_{0}}$ is also a non-trivial finite splitting with finite edge stabilizers because it is conjugate to the action $G_{0}\curvearrowright T_{e_{0}}$. For $e\notin G\cdot e_{0}$ we define $T_{e}$ to be a single vertex and we define $gT_{e}=T_{ge}$. Altogether this gives us an action of $G$ on $\sqcup_{e}T_{e}$ that is compatible with the action on $T$, and that satisfies properties (a) and (b) from Theorem 4.1.
4.4 The pocset $\mathcal{P}$
Define an equivalence relation on $V\mathfrak{h}_{0}$ where $x\sim y$ if $x$ and $y$ are not separated in $\mathfrak{h}_{0}$ by any $g\delta$ with $g\in G_{0}$. Equivalently, $x\sim y$ if there is no $D\in\mathcal{P}_{0}$ with $x\in D$ and $y\notin D$. Let $[x]$ denote the equivalence class of $x$, and let $[x]^{*}=VX-[x]$.
Since $T_{0}$ is the cubing of $\mathcal{P}_{0}$, and $\mathcal{P}_{0}$ forms a wallspace $(V\mathfrak{h}_{0},\mathcal{P}_{0})$, we have a map $\lambda:V\mathfrak{h}_{0}\to VT_{0}$ as in Definition 3. Note that the equivalence classes $[x]$ are just the non-empty fibers of $\lambda$.
@@ 4.4 The pocset $\mathcal{P}$
The subgraph of $X$ spanned by vertices that lie in $\mathfrak{h}_{0}-\hat{\mathfrak{h}}_{0}$ and not in halfspaces $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$ is the $p$-preimage of a vertex in $T$, so it is a connected subgraph (see Definition 2). As $\delta$ is contained in $\mathfrak{h}_{0}$ but avoids halfspaces $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$, there exists a finite connected subgraph $F\subseteq\mathfrak{h}_{0}$ that contains $\delta$ but avoids halfspaces $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$. Let $L$ be the length of the longest embedded path in $F$.
@@
Let $g\in G$ with $g\mathfrak{h}_{0}^{*}\subseteq\mathfrak{h}_{0}$. Let $[x]$ and $[y]$ be the $\sim$-classes with $g\mathfrak{h}_{0}^{*}\subseteq[x]$ and $\mathfrak{h}_{0}^{*}\subseteq g[y]$ (which exist by Lemma 4.13). Then $g[y]^{*}$ is contained in a bounded neighborhood of $[x]$.
@@
Observe that $g[y]^{*}$ is contained in the union of $g\mathfrak{h}_{0}^{*}$ and $g[z]$ for $\sim$-classes $[z]\neq[y]$. We already know that $g\mathfrak{h}_{0}^{*}\subseteq[x]$, so it remains to show that the sets $g[z]$ for $[z]\neq[y]$ are contained in a bounded neighborhood of $[x]$.
@@
Let $z\in V\mathfrak{h}_{0}$ with $[z]\neq[y]$. Suppose $gz\notin[x]$. Suppose $g_{0}\delta$ separates $[z]$ from $[y]$ in $\mathfrak{h}_{0}$, with $g_{0}\in G_{0}$. So $gg_{0}\delta$ separates $gz$ from $g[y]$ in $g\mathfrak{h}_{0}$. Since $\mathfrak{h}_{0}^{*}\subseteq g[y]$, we have that $gz\in\mathfrak{h}_{0}$ and that $gg_{0}\delta$ separates $gz$ from $\mathfrak{h}_{0}^{*}$ in $g\mathfrak{h}_{0}$. As $gz\notin[x]$, there is some $g_{1}\in G_{0}$ such that $g_{1}\delta$ separates $gz$ from $[x]$ in $\mathfrak{h}_{0}$. And as $g\mathfrak{h}_{0}^{*}\subseteq[x]$, we know that $g_{1}\delta$ also separates $gz$ from $g\mathfrak{h}_{0}^{*}$ in $\mathfrak{h}_{0}$.
Let $\gamma$ be a path in $X$ from $gz$ to $[x]$. We claim that $\gamma$ intersects the $L$-neighborhood of $gg_{0}\delta$. Suppose not. As $gg_{0}\delta$ separates $gz$ from $\mathfrak{h}_{0}^{*}$ in $g\mathfrak{h}_{0}$, and $g_{1}\delta$ separates $gz$ from $g\mathfrak{h}_{0}^{*}$ in $\mathfrak{h}_{0}$, we see that $\gamma$ will stay in $\mathfrak{h}_{0}\cap g\mathfrak{h}_{0}$ unless it intersects $g_{1}\delta$. Since $g_{1}\delta$ separates $gz$ from $[x]$ in $\mathfrak{h}_{0}$, we deduce that $\gamma$ must intersect $g_{1}\delta$. Let $z^{\prime}$ be the first point of intersection. Observe that $gg_{0}\delta$ separates $z^{\prime}$ from $\mathfrak{h}_{0}^{*}$ in $g\mathfrak{h}_{0}$. We also have $z^{\prime}\in g_{1}F$, so there is a path $\gamma^{\prime}$ in $g_{1}F$ from $z^{\prime}$ to $\hat{\mathfrak{h}}_{0}\subseteq\mathfrak{h}_{0}^{*}$ of length at most $L$. We have $g_{1}F\subseteq g\mathfrak{h}_{0}$ by construction of $F$, so $\gamma^{\prime}$ must cross $gg_{0}\delta$. Hence $\gamma$ intersects the $L$-neighborhood of $gg_{0}\delta$, as claimed.
@@
We have shown that either $d(gz,gg_{0}\delta)\leq L$ or the $L$-neighborhood of $gg_{0}\delta$ separates $gz$ from $[x]$ in $X$. Suppose for the moment we are in the latter case. As $X$ is one-ended, the $L$-neighborhood of $\delta$ only has one unbounded complementary component. Since $X$ is locally finite, there are only finitely many bounded complementary components, so there is a constant $K\geq L$ such that the bounded complementary components are contained in the $K$-neighborhood of $\delta$. Since $g\mathfrak{h}_{0}^{*}\subseteq[x]$, we know that $[x]$ is unbounded, so $gz$ is in a bounded complementary component of the $L$-neighborhood of $gg_{0}\delta$, and $d(gz,gg_{0}\delta)<K$. Thus in either case we get $d(gz,gg_{0}\delta)<K$. Then $d(gz,g\mathfrak{h}_{0}^{*})<K+L$, which in turn implies that $d(gz,[x])<K+L$. ∎
$gg_{0}\delta$$\mathfrak{h}_{0}$$g\mathfrak{h}_{0}$$g[y]$$gz$
$g_{1}\delta$$\mathfrak{h}_{0}$$g\mathfrak{h}_{0}$$[x]$$gz$$z^{\prime}$$\gamma$
@@
Figure 2: Cartoon picture for the proof of Lemma 4.14.

### v4

@@ 4.3 Halfspace cuts
$\mathfrak{h}$$\mathfrak{h}^{\prime}$CCDD
Figure 1: Cartoon picture for the proof of Lemma 4.10. Here $D$ is the vertex set of an unbounded component of the induced subgraph of $C-\mathfrak{h}^{\prime}$.
@@
The modification of $X$ we consider is where we fix an integer $n\geq 1$, and replace each edge in each wall of $X$ with $n$ edges that join the same pair of vertices. Let $X_{n}$ be the resulting graph, which is endowed with a collection of halfspaces and walls that are inherited from $X$ in the obvious way. Any halfspace cut $(\mathfrak{h},C)$ in $X$ is still a halfspace cut in $X_{n}$, and vice versa. The only difference is that the edge boundary $\delta C$ might become larger in $X_{n}$ because of the extra edges. More precisely, the numbers $W(C)$ and $|\delta C|$ both increase by (n1)W(C)(n-1)W(C) when passing from $X$ to $X_{n}$. Note that a halfspace cut is $W$-minimal with respect to $X$ if and only if it is $W$-minimal with respect to $X_{n}$, but the property of having minimal $|\delta C|$ may change when passing between $X$ and $X_{n}$. Let $w$ be the smallest value of $W(C)$ among halfspace cuts $(\mathfrak{h},C)$ in $X$, and let $N$ be the smallest value of $|\delta C|$ among $W$-minimal halfspace cuts $(\mathfrak{h},C)$ in $X$. Put $n=N+1$. Then the smallest possible value of $|\delta C|$ in $X_{n}$ is N+(n1)wN+(n-1)w, and this is only achieved by $W$-minimal halfspace cuts, as required. ∎
Now let $(\mathfrak{h}_{0},C)$ be a halfspace cut with $|\delta C|$ minimal. Lemmas 4.10 and 4.11 imply that each halfspace $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$ satisfies either $V\mathfrak{h}\subseteq C$ or $V\mathfrak{h}\subseteq V\mathfrak{h}_{0}-C$. Furthermore, by Theorem 4.6 we can assume that the pocset $\mathcal{P}(\mathfrak{h}_{0},C)$ has no pairs of transverse elements.
@@ 4.4 The trees $T_{e}$
For brevity we will write $G_{0}=G_{\mathfrak{h}_{0}}$ for the $G$-stabilizer of $\mathfrak{h}_{0}$ and we will write $\delta$ for the edge boundary $\delta C$. Note that $\mathfrak{h}_{0}$ corresponds to some edge $e_{0}\in ET$ (i.e. $p(\mathfrak{h}_{0})\subseteq T$ is bounded by $\hat{e}_{0}$) and $G_{0}=G_{e_{0}}$. As in Subsection 4.2, we can form the pocset $\mathcal{P}(\mathfrak{h}_{0},C)$. Now consider the subpocset
@@
We can then define the trees $T_{e}$ required for Theorem 4.1, and the action of $G$ on $\sqcup_{e}T_{e}$. Indeed, recall that there is an edge $e_{0}\in ET$ corresponding to $\mathfrak{h}_{0}$ (i.e. $p(\mathfrak{h}_{0})\subseteq T$ is bounded by $\hat{e}_{0}$). Put
@@
for each $i\in\Omega$. Note that the $G$-stabilizer of $T_{e_{0}}$ is $G_{0}=G_{e_{0}}$, and the action $G_{0}\curvearrowright T_{e_{0}}$ is a non-trivial finite splitting with finite edge stabilizers by Lemmas 4.12 and 4.13. Similarly, the $G$-stabilizer of $T_{g_{i}e_{0}}$ is $g_{i}G_{0}g_{i}^{-1}=G_{g_{i}e_{0}}$, and the action $G_{g_{i}e_{0}}\curvearrowright T_{g_{i}e_{0}}$ is also a non-trivial finite splitting with finite edge stabilizers because it is conjugate to the action $G_{0}\curvearrowright T_{e_{0}}$. For $e\notin G\cdot e_{0}$ we define $T_{e}$ to be a single vertex and we define $gT_{e}=T_{ge}$. Altogether this gives us an action of $G$ on $\sqcup_{e}T_{e}$ that is compatible with the action on $T$, and that satisfies properties (a) and (b) from Theorem 4.1.
4.5 An equivalence relation on $V\mathfrak{h}_{0}$
Define an equivalence relation on $V\mathfrak{h}_{0}$ where $x\sim y$ if $x$ and $y$ are not separated in $\mathfrak{h}_{0}$ by any $g\delta$ with $g\in G_{0}$. Equivalently, $x\sim y$ if there is no $D\in\mathcal{P}_{0}$ with $x\in D$ and $y\notin D$. Let $[x]$ denote the equivalence class of $x$, and let $x]^{*}=VX-[x]$.
Since $T_{0}$ is the cubing of $\mathcal{P}_{0}$, and $\mathcal{P}_{0}$ forms a wallspace $(V\mathfrak{h}_{0},\mathcal{P}_{0})$, we have a map $\lambda:V\mathfrak{h}_{0}\to VT_{0}$ as in Definition 3. Note that the equivalence classes $[x]$ are just the non-empty fibers of $\lambda$. We now prove two lemmas about how these equivalence classes interact with halfspaces $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$.
@@ 4.5 An equivalence relation on $V\mathfrak{h}_{0}$
Suppose $g\in G$ with $g\mathfrak{h}_{0}^{*}\subseteq\mathfrak{h}_{0}$. Then there exist $\sim$-classes $[x]$ and $[y]$ with $gV\mathfrak{h}_{0}^{*}\subseteq[x]$ and $V\mathfrak{h}_{0}^{*}\subseteq g[y]$. Furthermore, $g[y]^{*}$ is contained in a bounded neighborhood of $[x]$.
@@
Note that $g\mathfrak{h}_{0}^{*}\subsetneq\mathfrak{h}_{0}$ since $G$ acts on $T$ without edge inversions. The existence of $[x]$ with $gV\mathfrak{h}_{0}^{*}\subseteq[x]$ then follows immediately from Lemma 4.14. Meanwhile, $g\mathfrak{h}_{0}^{*}\subsetneq\mathfrak{h}_{0}$ implies $\mathfrak{h}_{0}^{*}\subsetneq g\mathfrak{h}_{0}$ by taking complementary halfspaces on both sides. Hence $g^{-1}\mathfrak{h}_{0}^{*}\subsetneq\mathfrak{h}_{0}$, and there exists $[y]$ with $g^{-1}V\mathfrak{h}_{0}^{*}\subset[y]$ by Lemma 4.14. Applying $g$ to both sides yields the desired inclusion $V\mathfrak{h}_{0}^{*}\subseteq g[y]$.
@@
It remains to prove the second part of the lemma. By definition, $y]^{*}$ is the union of $V\mathfrak{h}_{0}^{*}-V\mathfrak{h}_{0}$ and the sets $[z]$ for $\sim$-classes $z]\neq[y]$. Hence $g[y]^{*}$ is contained in the union of $g\mathfrak{h}_{0}^{*}$ and $g[z]$ for $\sim$-classes $z]\neq[y]$. We already know that $gV\mathfrak{h}_{0}^{*}\subseteq[x]$, so it remains to show that the sets $g[z]$ for $z]\neq[y]$ are contained in a bounded neighborhood of $[x]$.
@@
Let $z\in V\mathfrak{h}_{0}$ with $z]\neq[y]$. Suppose $gz\notin[x]$. Suppose $g_{0}\delta$ separates $[z]$ from $[y]$ in $\mathfrak{h}_{0}$, with $g_{0}\in G_{0}$. So $gg_{0}\delta$ separates $gz$ from $g[y]$ in $g\mathfrak{h}_{0}$. Since $V\mathfrak{h}_{0}^{*}\subseteq g[y]$, we have that $gz\in\mathfrak{h}_{0}$ and that $gg_{0}\delta$ separates $gz$ from $\mathfrak{h}_{0}^{*}$ in $g\mathfrak{h}_{0}$ (see the left-hand picture in Figure 2). On the other hand, as $gz\notin[x]$, there is some $g_{1}\in G_{0}$ such that $g_{1}\delta$ separates $gz$ from $[x]$ in $\mathfrak{h}_{0}$ (see the right-hand picture in Figure 2). And as $gV\mathfrak{h}_{0}^{*}\subseteq[x]$, we know that $g_{1}\delta$ also separates $gz$ from $g\mathfrak{h}_{0}^{*}$ in $\mathfrak{h}_{0}$.
The subgraph of $X$ spanned by vertices that lie in $\mathfrak{h}_{0}-\hat{\mathfrak{h}}_{0}$ and not in halfspaces $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$ is the $p$-preimage of a vertex in $T$, so it is a connected subgraph (see Definition 2). By construction of $(\mathfrak{h}_{0},C)$ at the end of Subsection 4.3, each edge in $\delta$ is contained in $\mathfrak{h}_{0}$ but not in any of the halfspaces $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$ (although it might have one endpoint in $\mathfrak{h}$). As a result, there exists a finite connected subgraph $F\subseteq\mathfrak{h}_{0}$ that contains $\delta$ and such that $F\subseteq\mathfrak{h}^{*}$ for any halfspace $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$. Let $L$ be the diameter of $F$. Note that $F$ and $L$ do not depend on the choice of $z$.
Let $\gamma$ be a path in $X$ from $gz$ to $[x]$. We claim that $\gamma$ intersects the $L$-neighborhood of $gg_{0}\delta$.
  1. (1)
    Suppose $\gamma$ intersects $\mathfrak{h}_{0}^{*}\cup g\mathfrak{h}_{0}^{*}$. Let $z^{\prime}$ be the first point of $\gamma$ that lies in $\mathfrak{h}_{0}^{*}\cup g\mathfrak{h}_{0}^{*}$. So the initial segment $\gamma^{\prime}\subseteq\gamma$ from $gz$ to $z^{\prime}$ lies in $\mathfrak{h}_{0}\cap g\mathfrak{h}_{0}$. As $gg_{0}\delta$ separates $gz$ from $\mathfrak{h}_{0}^{*}$ in $g\mathfrak{h}_{0}$, and $g_{1}\delta$ separates $gz$ from $g\mathfrak{h}_{0}^{*}$ in $\mathfrak{h}_{0}$, it must be that $\gamma^{\prime}$ intersects either $gg_{0}\delta$ or $g_{1}\delta$. If $\gamma^{\prime}$ intersects $gg_{0}\delta$ then the claim is proved, so suppose $\gamma^{\prime}$ intersects $g_{1}\delta$ but not $gg_{0}\delta$.
    Observe that $\gamma^{\prime}\cap g_{1}\delta\subseteq g_{1}F$. By Lemma 4.8, $g_{1}\delta$, and hence $g_{1}F$, intersects $\hat{\mathfrak{h}}_{0}$. By definition, $F\subseteq\mathfrak{h}^{*}$ for any halfspace $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$, so $g_{1}F$ has the same property; in particular $g_{1}F\subseteq g\mathfrak{h}_{0}$. As $gg_{0}\delta$ separates $gz$ from $\mathfrak{h}_{0}^{*}$ in $g\mathfrak{h}_{0}$, it must also separate $\gamma^{\prime}\cap g_{1}\delta$ from $\mathfrak{h}_{0}^{*}$ in $g\mathfrak{h}_{0}$. Since $g_{1}F$ is connected and lies in $g\mathfrak{h}_{0}$, we deduce that $gg_{0}\delta$ separates $\gamma^{\prime}\cap g_{1}\delta$ from $\mathfrak{h}_{0}^{*}$ in $g_{1}F$. As $L$ is the diameter of $F$, and hence $g_{1}F$, we see that $\gamma^{\prime}\cap g_{1}\delta$ lies in the $L$-neighborhood of $gg_{0}\delta$, as required.
  2. (2)
    Now suppose $\gamma$ does not intersect $\mathfrak{h}_{0}^{*}\cup g\mathfrak{h}_{0}^{*}$. Then $\gamma$ is contained in $\mathfrak{h}_{0}\cap g\mathfrak{h}_{0}$. As $g_{1}\delta$ separates $gz$ from $[x]$ in $\mathfrak{h}_{0}$, it must be that $\gamma$ intersects $g_{1}\delta$. If $\gamma$ also intersects $gg_{0}\delta$ then the claim is proved, so suppose $\gamma$ intersects $g_{1}\delta$ but not $gg_{0}\delta$. We can then argue as in the second half of part (1) (with $\gamma$ in place of $\gamma^{\prime}$) to conclude that $\gamma$ intersects the $L$-neighborhood of $gg_{0}\delta$. This completes the proof of the claim.
@@
We have shown that either $d(gz,gg_{0}\delta)\leq L$ or the $L$-neighborhood of $gg_{0}\delta$ separates $gz$ from $[x]$ in $X$. Suppose for the moment we are in the latter case. As $X$ is one-ended, the $L$-neighborhood of $\delta$ only has one unbounded complementary component. Since $X$ is locally finite, there are only finitely many bounded complementary components, so there is a constant $K\geq L$ such that the bounded complementary components are contained in the $K$-neighborhood of $\delta$. Since $gV\mathfrak{h}_{0}^{*}\subseteq[x]$, we know that $[x]$ is unbounded, so $gz$ is in a bounded complementary component of the $L$-neighborhood of $gg_{0}\delta$, and $d(gz,gg_{0}\delta)\leq K$. Thus in either case we get $d(gz,gg_{0}\delta)\leq K$. Now $gg_{0}F$ contains $gg_{0}\delta$ and intersects $g\mathfrak{h}_{0}^{*}$, and $gg_{0}F$ has diameter $L$, so $d(gz,g\mathfrak{h}_{0}^{*})\leq K+L$. In turn this implies $d(gz,[x])\leq K+L$. As $K$ and $L$ do not depend on the choice of $z$, this implies that all the sets $g[z]$ with $z\in V\mathfrak{h}_{0}$ and $z]\neq[y]$ are contained in a bounded neighborhood of $[x]$, as required. ∎
$gg_{0}\delta$$\mathfrak{h}_{0}$$g\mathfrak{h}_{0}$g[y]g[y]gzgz
[…]
Figure 2: Cartoon picture for the proof of Lemma 4.15. The two pictures appear mutually contradictory, and indeed the proof shows that this situation can only happen when $gz$ is uniformly close to both $gg_{0}\delta$ and $g_{1}\delta$, so in reality the picture would look more like Figure 3.
[…]
@@
Figure 3: Cartoon picture for the proof of Lemma 4.15. A depiction of how the situation might actually look.
4.6 The pocset $\mathcal{P}$